$SPAD/src/input parabola.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool parabola.output
)set message test on
)set message auto off
)clear all
--S 1 of 1
draw(curve(t**2 + 2*t - 1,t**2 + t - 2),t = -4..3)
--R
--R   Compiling function %B with type DoubleFloat -> DoubleFloat
--R   Compiling function %D with type DoubleFloat -> DoubleFloat
--R   Graph data being transmitted to the viewport manager...
--R   AXIOM2D data being transmitted to the viewport manager...
--R
--R   (1)  TwoDimensionalViewport: "t^2+2*t+-1"
--R                                                 Type: TwoDimensionalViewport
--E 1
)spool
)lisp (bye)

———-

References^{[1] nothing}

4